# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v5

@@ 3.1. Methodology for Simulating Missingness
Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Additional information on other datasets is available in Appendix B.2.
@@ 3.3.2. Model Performance
Broadly speaking, fairness metrics that are based on error rates are consistent with formal equality of opportunity, while those that are based on selection rates (SRD and DI) are consistent with substantive equality of opportunity (Khan et al., 2022). Which specific metric is more relevant depends on the application domain and on the stakeholder (Narayanan, 2018). SRD focuses on absolute disparities in selection rates (e.g., fixed quotas used in college admissions), while DI focuses on relative disparities (e.g., enforcing the 4/5th rule for non-discrimination in hiring in the U.S.). From the point of view of individuals being selected, TPRD is typically used when assessing fairness in opportunity allocation, such as hiring or loan approvals, where a positive outcome should be equally accessible across groups, while TNRD is most relevant in risk assessment and exclusion decisions, like medical diagnoses, where fairness in avoiding false negatives (e.g., failing to detect a disease) matters.
@@ 3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. The core component is the benchmark controller, which executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values using the selected MVI technique(s) and preprocesses the data with standard scaling for numerical features and one-hot encoding for categorical features. Next, the controller trains ML models, including hyperparameter tuning. The evaluation module then assesses imputation quality and model performance. For comprehensive profiling, it uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability, and fairness metrics across multiple sensitive attributes and their intersections.
@@ 4.1. Correctness of the Predictive Model
Note that the F1 score on the bank dataset is extremely low (0.32). This is due to severe class imbalance in the dataset (the base rate is 0.117, as reported in Table 9 in Appendix B). Further, from Figure 25 in Appendix C.2, we can see that the accuracy of the lgbm__\__clf is 0.89, which is marginally better than the accuracy reported using Logistic Regression (0.885) in (Le Quy et al., 2022) in Table 15. Also note that a model trained on imputed data can sometimes outperform one trained on clean data, as for german and heart on MCAR. We hypothesize that this happens when the model trained on cleaned data learns spurious correlations (e.g., due to noisy or erroneous values, especially in continuous features). However, when we simulate nulls, MVI techniques that leverage information from other rows could impute more accurate values, breaking some of these spurious correlations and resulting in slightly better performance.
@@ 4.3. Stability of the Predictive Model
Figure 24 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually. The reason is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
@@ 4.4. Imputation Quality and Fairness
Imputation Quality. We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)). We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
@@
Further, imputation fairness is insufficiently predictive of model fairness. For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, and median-mode and clustering have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02. Conversely, on diabetes under MAR, auto-ml, miss-forest, median-dummy and clustering all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12). We see similar trends for other imputation fairness metrics such as F1 score difference (Figure 7(a)) and RMSE difference (Figure 7(b)).
@@ 5.2. Fairness of the Predictive Model
Test set missingness. We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 53 in Appendix 52 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
@@ 6. Running Time
Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime. Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time. A particularly noteworthy comparison is between gain and edit-gain. As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models. Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 4. Importantly, with this running time acceleration, edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
@@ 7. Summary of Experimental Findings
Do not drop your nulls! Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our experiments show that deletion is the least effective strategy for model accuracy, fairness, and stability, especially in setups where each row contributes valuable information. Logically, deletion results in data loss, but its appropriateness depends on data quality rather than quantity. If rows are complete duplicates or contain erroneous values, deletion may be justified.
Multiple imputation shows mixed results.
There is conflicting evidence on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are similarly mixed and somewhat unexpected. Feng (2023) and Le Morvan et al. (2021) argue that MI outperforms impute-then-classify approaches in predictive performance, while Graham (2009) and McNeish (2017) find MI effective even with limited data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets and is less stable than simpler MVI techniques. This likely stems from the complexity-stability trade-off (Domingos, 2000; Breiman, 2001): MI employs a more complex model class that minimizes empirical loss but exhibits greater prediction variance under small training set perturbations.
Fairness is highly missingness-specific.
We find that no MVI technique is consistently fairness-preserving, corroborating the findings of  Zhang and Long (2021b) and Guha et al. (2024). Further, we find that fairness is highly sensitive to changes in missingness rates and missingness mechanisms between training and test sets, which are likely to occur in practice, and therefore a cause for ethical concern.
Imputation quality and fairness metrics are often insufficiently predictive of the correctness and fairness of downstream models.
Shadbahr et al. (2023) argue that distributional imputation quality metrics better predict model performance than discrepancy metrics. However, we find that neither reliably indicates downstream performance (F1), as strong learners can compensate for poor imputation quality. Moreover, imputation fairness metrics do not predict model fairness: perfect imputation fairness can still lead to model unfairness, while models trained with imputation-unfair MVI techniques can achieve good downstream fairness.
@@
Model stability depends more on the dataset size and MVI technique than on the missingness scenario.
@@
We find that for large datasets even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques preserve stability, while deletion, statistical imputation and more complex ML-based approaches all worsen stability.
Sensitivity to train and test missingness rates.
@@
Model performance (F1) is more affected by test missingness than training missingness. Fairness, however, is highly sensitive to both. Model stability is largely unaffected by test missingness, but most MVI techniques become more unstable with higher training missingness.
Existing MNAR-specific methods are insufficient.
@@
MNAR is theoretically the hardest setting to model. Empirically, we find that MVI techniques perform poorly with MNAR during testing, even the two existing MNAR-specific methods (not-miwae and mnar-pvae). For instance, experimental results show that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR. This performance gap arises because these methods rely on assumptions that often fail in real-world settings and are evaluated on limited, mostly large datasets with a single scenario, lacking the diversity of our benchmark. Developing improved MNAR-specific MVI methods and establishing more comprehensive benchmarks for their evaluation are critical directions for future research.
Deep learning outperforms tree-based models on large datasets.
@@
Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that Figures 22, 23, and 24 show only the best-performing models (by F1) for clarity. Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022). However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
Best performing approaches under single- and multi-mechanism missingness can be different.
@@
A key novelty of our study is the evaluation under mixed or multi-mechanism missingness—a more complex but realistic pattern likely encountered in practice. We find that performance trends vary between single- and multi-mechanism missingness. Figure 11 shows that model types like gandalf__\__clf and lgbm__\__clf are uncorrelated with performance metrics under single-mechanism missingness but become correlated under mixed missingness: F1 correlations are $\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf, while fairness (TPRD) correlations are $\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf.
It’s complicated!
@@
Prior studies (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) have found no universally best missing value imputation technique for predictive performance. Our findings reinforce this, revealing trade-offs between F1, fairness, and stability that depend on the predictive model’s architecture. For instance, under multi-mechanism missingness (Figure 11), logistic regression (lr__\__clf) correlates positively with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively with fairness (TPRD, $\rho$=-0.23). Random forest (rf__\__clf) shows weak positive correlations with fairness (TPRD, $\rho$=0.16) and F1 ($\rho$=0.25) but a negative correlation with stability ($\rho$=-0.24). Deep table-learning (gandalf__\__clf) is strongly correlated with fairness (TPRD, $\rho$=0.55), weakly with stability ($\rho$=0.12), and negatively with F1 ($\rho$=-0.24). This is not bad news, but rather an acknowledgement of the complexity of learning from incomplete data and the need for rigorous, holistic evaluation protocols, such as those demonstrated in this work, to select the most suitable imputation method and model architecture for a given task.
@@ 8. Conclusions, Limitations and Future Work
Limitations. We started with clean datasets, carefully designed meaningful missingness scenarios, and simulated them using error injectors. A limitation of our work is that we evaluate performance on synthetically generated — rather than on naturally occurring — missingness. This is a common limitation of work in this area: we often do not have a clean ground truth to compare against, and so we make the assumption that injecting synthetic errors will only exacerbate the effect of existing (unknown) errors. With this assumption, any degradation in performance we see is an underestimate of the possible degradation we can expect by comparing with an ideal clean ground truth.
Future work. Creating research datasets with naturally occurring missingness shifts is a promising direction. While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture. The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle. Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.

### v6

@@ 3.1. Methodology for Simulating Missingness
Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importance values in Figure 2. Additional information on other datasets is available in Appendix B.2.
@@ 3.3.2. Model Performance
Fairness metrics based on error rates align with formal equality of opportunity, while those based on selection rates (SRD, DI) reflect substantive equality (Khan et al., 2022). The choice of metric depends on the domain and stakeholder (Narayanan, 2018). SRD captures absolute disparities (e.g., fixed quotas in college admissions), while DI measures relative disparities (e.g., the 4/5th rule in U.S. hiring). From the individual’s perspective, TPRD suits opportunity allocation (e.g., hiring, loans), ensuring access to positive outcomes, whereas TNRD applies to exclusion decisions (e.g., medical diagnoses), emphasizing fairness in avoiding false negatives.
@@ 3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. Its core component, the benchmark controller, executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values using selected MVI technique(s) and preprocesses data via standard scaling (numerical) and one-hot encoding (categorical), and then trains ML models with hyperparameter tuning. The evaluation module assesses imputation quality and model performance. For comprehensive profiling, it uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability, and fairness metrics across multiple sensitive attributes and their intersections.
@@ 4.1. Correctness of the Predictive Model
The F1 score on the bank dataset is low (0.32), due to severe class imbalance (base rate 0.117, see Table 9 in Appendix B). Figure 25 in Appendix C.2 shows that lgbm__\__clf achieves 0.89 accuracy, slightly outperforming Logistic Regression (0.885) reported in (Le Quy et al., 2022) in Table 15. Interestingly, models trained on imputed data can sometimes outperform those trained on clean data, as seen for german and heart under MCAR. We hypothesize this occurs when models trained on cleaned data learns spurious correlations (e.g., from noisy or erroneous values), while MVI methods may impute more accurate values, mitigating such artifacts and improving performance.
@@ 4.3. Stability of the Predictive Model
Figure 24 shows label stability of models trained with different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends primarily on dataset characteristics (especially size) and model type. Impute-then-classify methods like miss-forest, auto-ml, and datawig, which perform best on F1, also match the stability of models trained on clean data across missingness types and dataset sizes. nomi, which performs best on accuracy and training time, shows comparable stability to these top methods across datasets. clustering, which performed poorly on F1, is likewise unstable on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable exception is german under mixed missingness, where clustering is competitive despite underperforming on MCAR, MAR, or MNAR individually. This may be because imputation accuracy affects data uncertainty, which in turn drives model uncertainty (Gal et al., 2016).
@@ 4.4. Imputation Quality and Fairness
Imputation Quality. MVI techniques with widely varying imputation quality can yield models with similar F1, suggesting that imputation correctness is not a reliable predictor of downstream performance (Zhang and Long, 2021a). For instance, in Figure 6(a) on diabetes, median-dummy has imputation F1 near 0; auto-ml, miss-forest, and nomi are near 1; others fall between 0.5–0.6, yet all produce models with F1 close to 1. This pattern holds across datasets, missingness types, and for numerical columns (Figure 6(b)). Similar trends are observed for KL divergence (Figure 6(c)), reinforcing that neither discrepancy-based nor distributional metrics reliably predict downstream model performance, contradicting Shadbahr et al. (2023)’s claim.
@@
Further, imputation fairness is insufficiently predictive of model fairness. For example, on german under mixed missingness, median-dummy has KL difference close to 0, datawig and miss-forest have KL difference between 1 and 1.5, clustering and median-mode have KL difference between -1 and -1.5, but the models trained using all five of these techniques have a TPRD close to -0.02. Conversely, on diabetes under MAR, auto-ml, clustering, miss-forest, and median-dummy all have near perfect imputation fairness (KL difference close to 0), but different model fairness (TPRD between 0.04 and 0.12). We see similar trends for other imputation fairness metrics such as F1 score difference (Figure 7(a)) and RMSE difference (Figure 7(b)).
@@ 5.2. Fairness of the Predictive Model
Test set missingness. We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to such shifts. Figure 53 in Appendix 52 shows the TPRD of the Random Forest model on diabetes as test missingness increases. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) as test error rate rises, though not always monotonically. miss-forest, auto-ml, and nomi are robust to increases in test missingness rate under all scenarios. Simpler methods such as deletion, clustering, median-mode and median-dummy show no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example, with deletion and clustering, TPRD increases with test missingness in scenarios S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but decreases in S2 (MAR train, MAR test).
@@ 6. Running Time
Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient. Notably, nomi delivers accuracy on par with miss-forest, datawig, and auto-ml, successfully balancing imputation accuracy and training time. A key comparison is between gain and edit-gain. As explained in Appendix A.1, edit-gain achieves a 28x speedup on folk_emp, with even greater improvements for smaller datasets, as shown in Table 4, while maintaining comparable accuracy to gain , as shown in Figures 18 and 19.
@@ 7. Summary of Experimental Findings
Do not drop your nulls! Building on prior evidence (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we confirm that deletion is the least effective strategy for model accuracy, fairness, and stability—especially when each row holds valuable information. While deletion leads to data loss by design, its suitability depends more on data quality than quantity. If rows are duplicates or contain errors, deletion may be warranted.
Multiple imputation shows mixed results. There is conflicting evidence on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are similarly mixed and somewhat unexpected. Feng (2023) and Le Morvan et al. (2021) argue that MI outperforms impute-then-classify approaches in predictive performance, while Graham (2009) and McNeish (2017) find MI effective even with limited data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets and is less stable than simpler MVI techniques. This likely stems from the complexity-stability trade-off (Domingos, 2000; Breiman, 2001): MI employs a more complex model class that minimizes empirical loss but exhibits greater prediction variance under small training set perturbations.
Fairness is highly missingness-specific. We find that no MVI technique is consistently fairness-preserving, corroborating the findings of  Zhang and Long (2021b) and Guha et al. (2024). Further, we find that fairness is highly sensitive to changes in missingness rates and missingness mechanisms between training and test sets, which are likely to occur in practice, and therefore a cause for ethical concern.
Imputation quality and fairness metrics often fail to predict the correctness or fairness of downstream models. Shadbahr et al. (2023) argue that distributional imputation quality metrics better predict model performance than discrepancy metrics. However, we find that neither reliably predicts downstream performance, as strong learners can compensate for poor imputations. Moreover, imputation fairness does not predict model fairness: fair imputers can still yield unfair models, while models trained with fairness-poor MVI techniques can achieve good downstream fairness.
@@
Model stability depends more on the dataset size and MVI technique than on the missingness scenario. We find that for large datasets even simple statistical imputers can preserve stability. In contrast, for small datasets, only a few MVI techniques do so, while deletion, statistical imputation and complex ML-based MVI all worsen stability.
@@
Sensitivity to train and test missingness rates. Model performance (F1) is more affected by test missingness than training missingness. Fairness, however, is highly sensitive to both. Model stability is largely unaffected by test missingness, but most MVI techniques become more unstable with higher training missingness.
@@
Existing MNAR-specific methods are insufficient. MNAR is theoretically the hardest setting to model. MVI techniques, including MNAR-specific not-miwae and mnar-pvae, perform poorly under MNAR. For example, F1 and stability are more sensitive to missingness rates under MNAR than under MCAR or MAR. This performance gap stems from unrealistic assumptions and limited evaluations, typically on large datasets and a single scenario, lacking the diversity of our benchmark. Advancing MNAR-specific MVI methods and building comprehensive benchmarks are key future directions.
@@
Deep learning outperforms tree-based models on large datasets. Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that Figures 22, 23, and 24 show only the best-performing models (by F1) for clarity. Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022). However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
@@
Best-performing approaches under single- and multi-mechanism missingness can differ. A key contribution of our study is the evaluation under mixed (multi-mechanism) missingness — a complex pattern likely to arise in practice. We find that performance trends differ between single- and multi-mechanism settings. Figure 11 shows that model types like gandalf_clf and lgbm_clf are uncorrelated with performance metrics under single-mechanism missingness, but become correlated under mixed missingness: F1 correlations are $\rho=-0.24$ for gandalf_clf and $\rho=-0.37$ for lgbm_clf, while fairness (TPRD) correlations are $\rho=0.55$ and $\rho=-0.4$, respectively.
@@
It’s complicated! Prior studies (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) have found no universally best missing value imputation technique for predictive performance. Our findings reinforce this, revealing trade-offs between F1, fairness, and stability that depend on the predictive model’s architecture. For instance, under multi-mechanism missingness (Figure 11), logistic regression (lr__\__clf) correlates positively with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively with fairness (TPRD, $\rho$=-0.23). Random forest (rf__\__clf) shows weak positive correlations with fairness (TPRD, $\rho$=0.16) and F1 ($\rho$=0.25) but a negative correlation with stability ($\rho$=-0.24). Deep table-learning (gandalf__\__clf) is strongly correlated with fairness (TPRD, $\rho$=0.55), weakly with stability ($\rho$=0.12), and negatively with F1 ($\rho$=-0.24). This is not bad news, but a recognition of the complexity of learning from incomplete data, and of the need for rigorous, holistic evaluation protocols, like those used here, to identify the best imputation method and model architecture for a given task.
@@ 8. Conclusions, Limitations and Future Work
Limitations.
We began with clean datasets, designed meaningful missingness scenarios, and simulated them using error injectors. A key limitation is that we evaluate performance on synthetically generated—rather than naturally occurring—missingness. This is common in the field: clean ground truth is rarely available, so we assume that injecting synthetic errors only worsens the effects of existing (unknown) ones. Under this assumption, any observed performance degradation likely underestimates the true degradation relative to an ideal ground truth.
Future work.
Creating research datasets with naturally occurring missingness shifts is a promising direction. While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture. The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle. Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.